OpenStudy (mary1999):

enter an ordered pair corresponding to a 90 degree counterclockwise rotation of the point (14,10)about the origin followed by a reflection in the line x=6

OpenStudy (blockcolder):

When you rotate the point (a,b) about the origin counterclockwise at an angle of \(\theta \), then it gets mapped to the point \((a \cos\theta-b\sin\theta, a\sin\theta+y\cos\theta)\). The point (a, b), when reflected about a vertical line x=c, becomes (2c-a, b). From these two facts, you'll be able to answer your question.